# Unveiling Environmental Impacts of Large Language Model Serving: A Functional Unit View

### v1

@@ Abstract
Large language models (LLMs) offer powerful capabilities but come with significant environmental costs, particularly in carbon emissions. Existing studies benchmark these emissions but lack a standardized basis for comparison across models. To address this, we introduce the concept of a functional unit (FU) and develop FUEL, the first FU-based framework for evaluating LLM serving’s environmental impact. Through case studies on model size, quantization, and hardware, we uncover key trade-offs in sustainability. Our findings highlight the potential for reducing carbon emissions by optimizing model selection, deployment strategies, and hardware choices, paving the way for more sustainable AI infrastructure.
@@ A Functional Unit View
Yanran Wu Purdue University wu2187@purdue.edu                        Inez Hua Purdue University hua@purdue.edu                        Yi Ding Purdue University yiding@purdue.edu
@@ 1 Introduction
Large language models (LLMs) have been widely adopted in various industries due to their ability to perform complex language tasks Vu et al. (2024); Shen et al. (2024); Liu et al. (2024c). However, LLM serving comes with significant environmental impacts, particularly in terms of carbon emissions. For instance, processing a single prompt on ChatGPT produces over 4 grams of CO2eq Wong (2023), which is over 20x the carbon emissions generated by a web search query Griffiths (2020).
Recent studies have benchmarked the carbon emissions of LLM serving by analyzing performance (e.g., throughput, latency) and energy consumption, then modeling carbon emissions under varying conditions such as request rate, and input/output length Nguyen et al. (2024); Li et al. (2024d); Shi et al. (2024); Li et al. (2024b). However, these efforts have two limitations: (1) they focus on individual LLMs rather than cross-model comparisons, and (2) they lack a standardized basis for fair carbon emission comparisons. These gaps limit the broader applicability and fairness of their analyses.
Building on principles from life cycle assessment in environmental sustainability Klöpffer and Grahl (2014), we address these two limitations by introducing the concept of a functional unit (FU) as a standardized basis for comparing LLMs. In LLM serving, an FU represents a token generation defined by workload intensity, performance, and quality constraints. Using this, we develop FUEL, a Functional Unit-based Evaluation framework for evaluating the environment impact of LLMs. To demonstrate its effectiveness and generalizability, we conduct three case studies exploring model size, quantization, and hardware. Our key insights for building sustainable LLM serving systems include:
@@
  • Hardware: Newer hardware offers better performance but is not always greener due to higher embodied carbon. Older hardware can lower carbon emissions while meeting quality and performance constraints.
@@
  • Introducing and defining FU for LLM serving from environmental sustainability.
  • Developing FUEL, the first FU-based framework for assessing the environmental impact of LLM serving.
  • Conducting case studies on model size, quantization, and hardware impact on carbon emissions.
@@ 2 Related Work
LLM serving optimization. Prior work on LLM serving has primarily focused on optimizing performance and energy efficiency. Performance improvements can be categorized into model-level and system-level techniques. Model-side optimizations include quantization Lin et al. (2024); Frantar et al. (2022), sparsification Frantar and Alistarh (2023), and speculative decoding Leviathan et al. (2023). System-side approaches involve memory management Kwon et al. (2023), batching Agrawal et al. (2024); Yu et al. (2022), and kernel optimizations Dao et al. (2022). Additionally, efforts to enhance energy efficiency include solutions like Splitwise Patel et al. (2024b) and DynamoLLM Stojkovic et al. (2024). However, they have largely overlooked quality constraints.
@@ 3.2 Step 2: Define Functional Unit
Quality constraint. Quantitatively assessing output quality is challenging. While prior works (Zhong et al. (2022); Yuan et al. (2021); Jiang et al. (2023)) have introduced various methods, they depend on either specific datasets or the need for reference answers. After evaluating multiple quality metrics, we adopt the reward model Liu et al. (2024a), a common approach in reinforcement learning from human feedback training Ouyang et al. (2022). Our experiments show that the reward model’s scores align most closely with human preferences and effectively differentiate outputs across models. Using the reward model’s score, we define Qscore as a measure of output quality, where a higher Qscore reflects better quality and indicates that the output meets a certain quality threshold.
@@ 3.4 Step 4: Carbon Modeling
  • Operational carbon is calculated as the product of the energy consumed, $E_{\rm op}$, and the carbon intensity of the energy source (CI). Carbon intensity is defined as the amount of $CO_{\rm 2eq}$ emitted per kilowatt-hour ($kWh$) of electricity used Maji et al. (2022); Li et al. (2024a). The operational carbon emission is thus given by:
@@ 4.2 Evaluation Results
For Qwen, at a low Qscore of -5, smaller models emit less carbon. However, as Qscore increases, carbon emissions increase for all model sizes, with smaller models increasing at a faster rate. When Qscore exceeds 5, the smallest 7B model becomes the highest emitter. At Qscore 15, the 32B model has the lowest emissions, while the 7B model emits over 1.8× more. A similar trend is seen in Llama, where larger models become greener as quality requirement rise. We confirm that larger models produce higher-quality outputs with higher Qscores in Figure 21 in Section C.1. This underscores the need to balance model size and output quality for lower environmental impact.
@@
When are larger models greener? To examine when larger models become greener, we set FUs with a broader QPS range and two quality requirements: low (Qscore = -5) and high (Qscore = 15 for Qwen, 10 for Llama). Figure 3 shows that for Qwen, larger models (14B and 32B) save more carbon compared to the 7B model under high Qscore, with the 32B saving over 40%. However, under a low-quality requirement (Qscore = -5), larger models offer no advantage. A similar trend is seen for Llama, where the 13B model saves over 20% carbon compared to the 7B model at high quality. Thus, larger models become greener when output quality requirements are high.
@@ 5.2 Evaluation Results
To understand why AWQ does not always outperform FP16 in carbon savings, we analyze its impact on TTFT and TPOT speedup. Figures 9 and 10 show that TPOT sees some speedup at low QPS but slows down at high QPS, while TTFT is always slower than FP16. This is because quantization reduces weight size, but weights are dequantized back to 16-bit during inference, adding overhead. AWQ improves TPOT in memory-bound cases at low QPS by reducing memory transfer, but this advantage diminishes as QPS increases and computation grows. Since TTFT is compute-intensive, AWQ provides no speedup.
@@
To understand why W8A8 always outperforms FP16 in carbon savings, we analyze its impact on TTFT and TPOT speedup. Figures 13 and 14 show that W8A8 consistently speeds up TPOT and TTFT across all QPS ranges. This improvement comes from reducing both weight and activation precision, which decreases the amount of data movement and computation during inference. This makes W8A8 a more sustainable choice for LLM serving, as it strikes a balance between quality and performance.
@@ 6.2 Evaluation Results
The main advantage of advanced hardware like the H100 is its ability to produce higher-quality outputs and meet performance constraints, as shown in Figure 18. Although advanced hardware may not be greener, it provides better performance and supports higher-quality LLM serving at scale.
@@
Takeaway 4: Advanced hardware offers higher performance but is not always greener due to higher embodied carbon. Older hardware can achieve lower carbon emissions while still meeting quality and performance constraints.
@@ 7 Conclusion
This research aims to contribute to the development of sustainable and carbon efficient LLM serving systems. We are committed to conducting our work in a responsible manner, adhering to ethical guidelines and best practices. Our focus is on minimizing the environmental impact of LLM deployments while ensuring that the quality of the models and the performance of the systems meet the necessary standards for practical use.
We recognize the potential environmental consequences of the widespread use of LLMs, including energy consumption, electronic waste, and the environmental impact of hardware manufacturing. Therefore, we emphasize the importance of optimizing LLMs for lower energy and carbon emissions, not only in terms of performance but also through hardware reuse and longevity, as part of a more sustainable approach to AI infrastructure.
We strive to be transparent in our research methodologies and encourage further exploration of green AI practices. As we explore new avenues for improving LLM efficiency, we remain mindful of the broader social, economic, and environmental implications of deploying large-scale AI systems and aim to promote solutions that benefit both the technology and society at large.
@@
Appendix A Code Availability
Please find our code repository at https://anonymous.4open.science/r/FUEL-76EF. We provide a README file that offers instructions on how to get started and utilize FUEL.
@@ Appendix B Embodied Carbon Modeling
The packaging carbon emission $C_{\rm p}$ is calculated by the number of integrated circuits ($N_{\rm IC}$) with a packaging footprint. Following ACT, we use an average packaging overhead of 150 gCO2 per IC.
@@
In cloud environments or HPC clusters, it is often challenging to obtain details of DRAM specifications. Previous studies Li et al. (2023); Köhler et al. (2023) generally assume that the embodied carbon of DRAM is proportional to its capacity. Following prior work, we adopt a fixed rate of 65gCO2/GB to estimate the embodied carbon of DRAM.
@@ Appendix E Additional Results for Hardware Case Study
Figure 38 compares model and hardware combinations, further confirming our previous conclusion: older hardware can achieve lower carbon emissions. As shown in the figure, whether for the Qwen or Llama model families, the greenest choice at low QPS is consistently the L40 server. Once the hardware is fixed, we can apply insights from the model size case study to select the model size based on the quality requirement. This pattern reaffirms that choosing the optimal model and hardware combination requires a balance between performance needs and carbon efficiency.

### v2

# A Functional Unit View

@@ Abstract
Large language models (LLMs) offer powerful capabilities but come with significant environmental impact, particularly in carbon emissions. Existing studies benchmark carbon emissions but lack a standardized basis for comparison across different model configurations. To address this, we introduce the concept of functional unit (FU) as a standardized basis and develop FUEL, the first FU-based framework for evaluating LLM serving’s environmental impact. Through three case studies, we uncover key insights and trade-offs in reducing carbon emissions by optimizing model size, quantization strategy, and hardware choice, paving the way for more sustainable LLM serving. The code is available at https://github.com/jojacola/FUEL.
[…]
@@ A Functional Unit View
Yanran Wu, Inez Hua, Yi Ding Purdue University {wu2187, hua, yiding}@purdue.edu
@@ 1 Introduction
Large language models (LLMs) have been widely adopted in various industries due to their ability to perform complex language tasks Vu et al. (2024); Shen et al. (2024); Liu et al. (2024c). However, LLM serving comes with significant environmental impact, particularly in terms of carbon emissions. For instance, processing a single prompt on ChatGPT produces over 4 grams of CO2eq Wong (2023), which is over 20× the carbon emissions generated by a web search query Griffiths (2020).
Recent studies have benchmarked the carbon emissions of LLM serving by analyzing performance (e.g., throughput, latency) and energy consumption, then modeling carbon emissions under varying conditions such as request rate, and input/output length Nguyen et al. (2024); Li et al. (2024d); Shi et al. (2024); Li et al. (2024b). However, these efforts have two limitations: (1) they focus on individual LLM rather than cross-model comparisons, and (2) they lack a standardized basis for carbon emission comparisons. These gaps limit the broader applicability and fairness of their analyses.
Building on principles from life cycle assessment in environmental sustainability Klöpffer and Grahl (2014), we address these two limitations by introducing the concept of functional unit (FU) as a standardized basis for comparing LLMs. In LLM serving, an FU represents a token generation defined by workload intensity, performance, and quality constraints. Using this, we develop FUEL, a Functional Unit-based Evaluation framework for evaluating the environment impact of LLMs. To demonstrate its effectiveness and generalizability, we conduct three case studies exploring model size, quantization strategy, and hardware choice. Our key insights for building sustainable LLM serving systems include:
@@
  • Hardware: Newer hardware offers better performance but is not always greener due to higher embodied carbon. Older hardware can lower carbon emissions while meeting quality and performance constraints under certain conditions.
@@
  • Introducing FU as a standardized basis for comparing different LLMs.
  • Designing and implementing FUEL, the first FU-based framework for assessing the environmental impact of LLM serving.
  • Conducting three case studies on model size, quantization, and hardware to draw insights in building sustainable LLM serving systems.
@@ 2 Related Work
LLM serving optimization. Prior work on LLM serving has primarily focused on optimizing performance and energy efficiency. Performance improvements can be categorized into model-level and system-level techniques. Model-side optimizations include quantization Lin et al. (2024); Frantar et al. (2022), sparsification Frantar and Alistarh (2023), and speculative decoding Leviathan et al. (2023). System-side approaches involve memory management Kwon et al. (2023), batching Agrawal et al. (2024); Yu et al. (2022), and kernel optimizations Dao et al. (2022). Additionally, efforts to enhance energy efficiency include solutions like Splitwise Patel et al. (2024b) and DynamoLLM Stojkovic et al. (2024). However, they have largely overlooked quality constraints when considering performance and energy efficiency.
@@ 3.2 Step 2: Define Functional Unit
Quality constraint. Quantitatively assessing output quality is challenging. While prior works (Zhong et al. (2022); Yuan et al. (2021); Jiang et al. (2023)) have introduced various methods, they depend on either specific datasets or the need for reference answers. After evaluating multiple quality metrics, we adopt the reward model as a consistent quality evaluator for the model outputs across our experiments. The reward model is a common approach in reinforcement learning from human feedback (RLHF) training Ouyang et al. (2022). Specifically, we use the open-source pre-trained Skywork reward model Liu et al. (2024a), which ranks highly on the RewardBench benchmark Lambert et al. (2024), reflecting its strong performance on diverse language generation tasks. Our experiments show that its reward scores align well with human preferences and effectively differentiate output quality across models. Using the reward model’s score, we define Qscore as a measure of output quality, where a higher Qscore reflects better quality and indicates that the output meets a certain quality threshold.
@@ 3.4 Step 4: Carbon Modeling
  • Operational carbon is calculated as the product of the energy consumed, $E_{\rm op}$, and the carbon intensity of the energy source (CI). Carbon intensity is the amount of $CO_{\rm 2eq}$ emitted per kilowatt-hour ($kWh$) of electricity used Maji et al. (2022); Li et al. (2024a); Yan et al. (2025). The operational carbon emission is thus given by:
@@ 4.2 Evaluation Results
For Qwen, at a low Qscore of -5, smaller models emit less carbon. However, as Qscore increases, carbon emissions increase for all model sizes, with smaller models increasing at a faster rate. When Qscore exceeds 5, the smallest 7B model becomes the highest emitter. At Qscore 15, the 32B model has the lowest emissions, while the 7B model emits over 1.8× more. A similar trend is seen in Llama, where larger models become greener as quality requirement rise. We confirm that larger models produce higher-quality outputs with higher Qscores in Figure 3. This underscores the need to balance model size and output quality for lower carbon emissions.
These results demonstrate how FUEL provides a quantitative approach to assess the carbon efficiency of different models under consistent quality and performance requirements. In high-quality regimes, smaller models tend to emit more carbon per FU because fewer of their responses satisfy the criteria to be counted as functional units, leading to a higher average emission per FU.
@@
When are larger models greener? To examine when larger models become greener, we set FUs with a broader QPS range and two quality requirements: low (Qscore = -5) and high (Qscore = 15 for Qwen, 10 for Llama). Figure 4 shows that for Qwen, larger models (14B and 32B) save more carbon compared to the 7B model under high Qscore, with the 32B saving over 40%. However, under a low-quality requirement (Qscore = -5), larger models offer no advantage. A similar trend is seen for Llama, where the 13B model saves over 20% carbon compared to the 7B model at high quality, as shown in Figure 5. Thus, larger models become greener when output quality requirements are high.
@@ 5.2 Evaluation Results
To understand why AWQ does not always outperform FP16 in carbon savings, we analyze its impact on TTFT and TPOT speedup. Figures 10 and 11 show that TPOT sees some speedup at low QPS but slows down at high QPS, while TTFT is always slower than FP16. This is because quantization reduces weight size, but weights are dequantized back to 16-bit during inference, adding overhead. AWQ improves TPOT in memory-bound cases at low QPS by reducing memory transfer, but this advantage diminishes as QPS increases and computation grows. Since TTFT is compute-intensive, AWQ provides no speedup.
@@
To understand why W8A8 always outperforms FP16 in carbon savings, we analyze its impact on TTFT and TPOT speedup. Figures 14 and 15 show that W8A8 consistently speeds up TPOT and TTFT across all QPS ranges. This improvement comes from reducing both weight and activation precision, which decreases the amount of data movement and computation during inference. This makes W8A8 a more sustainable choice for LLM serving, as it strikes a balance between quality and performance.
@@ 6.2 Evaluation Results
The main advantage of advanced hardware such as the H100 lies in its higher computational throughput, memory capacity, and improved model support. These capabilities enable it to produce higher-quality outputs while satisfying latency and throughput constraints, as shown in Figure 19. However, this comes at the cost of significantly higher power consumption and embodied carbon. As a result, advanced hardware does not always lead to lower emissions per FU, especially when older hardware is sufficient to meet the same quality and performance requirements.
@@
Takeaway 4: Advanced hardware offers higher performance but is not always greener due to higher embodied carbon and power consumption. When quality and performance constraints can be met, older hardware can achieve lower carbon emissions per functional unit.
@@ Acknowledgments
This work contributes to the development of carbon-efficient LLM serving systems. We are committed to conducting our work in a responsible manner, adhering to ethical guidelines and best practices.
We recognize the potential environmental impact of the widespread use of LLMs, including energy consumption, electronic waste, and the environmental impact of hardware manufacturing. Therefore, we emphasize the importance of optimizing LLMs for lower energy and carbon emissions, not only in terms of performance but also through hardware reuse and longevity, as part of a more sustainable approach to AI infrastructure.
We are transparent in our research methodologies. As we explore new avenues for improving LLM efficiency, we remain mindful of the broader social, economic, and environmental implications of deploying large-scale AI systems and aim to promote solutions that benefit both the technology and society at large.
@@
Appendix A Appendix Overview
@@ Appendix B Embodied Carbon Modeling
The packaging carbon emission $C_{\rm p}$ is calculated by the number of integrated circuits ($N_{\rm IC}$) with a packaging footprint. Following ACT, we use an average packaging overhead of 150 gCO2eq per IC.
@@
In cloud environments or HPC clusters, it is often challenging to obtain details of DRAM specifications. Previous studies Li et al. (2023); Köhler et al. (2023) generally assume that the embodied carbon of DRAM is proportional to its capacity. Following prior work, we adopt a fixed rate of 65 gCO2eq/GB to estimate the embodied carbon of DRAM.
@@ Appendix E Additional Results for Hardware Case Study
Figure 38 compares model and hardware combinations, further confirming our previous conclusion: older hardware can achieve lower carbon emissions under certain conditions. As shown in the figure, whether for the Qwen or Llama model families, the greenest choice at low QPS is consistently the L40 server. Once the hardware is fixed, we can apply insights from the model size case study to select the model size based on the quality requirement. This pattern reaffirms that choosing the optimal model and hardware combination requires a balance between performance needs and carbon efficiency.